OpenStudy (anonymous):

Which is the solution of -5+2x^2=-6x

OpenStudy (anonymous):

First of all, try getting all terms on one side of the equation.

OpenStudy (anonymous):

\[2+ or -\sqrt{4-(-4)(6)(-5)} \over 5\]

OpenStudy (anonymous):

2x^2 - 5 = -6x 2x^2 + 6x -5 = 0 (2x

OpenStudy (anonymous):

that's as far as i got, i have to go though. See if you can figure it out. i don't think it's set up the way you put it

OpenStudy (anonymous):

the equation i put is one of the answers

OpenStudy (anonymous):

\[B. -5+or - \sqrt{25 - (4)(2)(6) }\]

OpenStudy (johnweldon1993):

So get everything on one side and set it equal to 0 (we'll use the quadratic formula here -5+2x^2=-6x becomes 2x^2 + 6x - 5 So we have everything we need... A = 2 B = 6 C = -5 In the quadratic formula \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Just plug in all your numbers and you'll have your answer

OpenStudy (anonymous):

Thank you @johnweldon1993

OpenStudy (johnweldon1993):

Anytime :)